# THE SIEGEL MODULAR GROUP IS THE LATTICE OF

### v2

@@ 1 Introduction
By the celebrated arithmeticity theorem of Margulis (see e.g. [19, Thm. 16.3.1]), any irreducible lattice $\Lambda$ in a semisimple Lie group $G$ of real rank greater than $1$ is also an arithmetic subgroup of $G$. In particular, if such a Lie group $G$ is a simple, any lattice is arithmetic. That is, there exists a ${\bf Q}$-embedding $\iota:G\rightarrow\mathrm{GL}_{n}({\bf R})$ such that $\Lambda$ is commensurable to the group $G({\bf Z})=\iota^{-1}(\iota(G)\cap\mathrm{GL}_{n}({\bf Z}))$ of integer points of $G$. For our purposes, the utility of this identification of lattices with arithmetic subgroups is that arithmetic subgroups are, in some sense, easier to describe. However, as any individual arithmetic subgroup of $G$ is associated to a corresponding embedding of $G$ into $\mathrm{GL}_{n}({\bf R})$, investigating all arithmetic subgroups involves the abstraction of treating all ${\bf Q}$-isomorphic copies of $G$ in $\mathrm{GL}_{n}({\bf R})$ equally.
We therefore let $\mathbb{G}$ be an algebraic group over a number field $K$, for which there exists an embedding $\upsilon_{0}:K\rightarrow{\bf R}$ which induces an ${\bf R}$-isomorphism from $\mathrm{Sp}_{2n}$ (the standard split form of the symplectic group) to $\mathbb{G}$. This is enough to capture any candidate for $\Gamma$ that we will want to consider. Moreover, it leads to a useful alternative characterization of $\Gamma$: Rohlfs gave a cohomological criterion [26, Satz 3.5] (see also [5, Prop. 1.4]) for maximality of arithmetic subgroups, according to which the lattice $\Gamma\subseteq\mathrm{Sp}_{2n}({\bf R})$ must be conjugate to the normalizer of a so-called principal arithmetic subgroup $\Lambda$ of $\mathbb{G}(K)$. Such subgroups are distinguished by being topologically well-behaved in relation to the group $\mathbb{G}(\mathbb{A}_{K})$ of adelic points of $\mathbb{G}$. In particular, $\Lambda$ is given by the intersection $\mathbb{G}(K)\cap\Pi_{\upsilon<\infty}P_{\upsilon}$ for a certain (topologically coherent) family of compact-open (more precisely, parahoric) subgroups $P_{\upsilon}\subseteq\mathbb{G}(K_{\upsilon})$, where $\upsilon$ ranges over all finite places of $K$. These facts make the problem of computing the covolume of $\Lambda$, and thereby identifying a lattice $\Gamma$ of minimal covolume, amenable to a wide range of analytic and algebraic tools.
@@ 1.2 Notation and conventions
  • $\Lambda$ will denote the principal arithmetic subgroup of $\mathbb{G}_{\infty}$ defined by the collection of parahorics $\mathopen{}\mathclose{{}\left\{P_{\upsilon}:\upsilon<\infty}\right\}$, for which $\Gamma$ is its normalizer in the group $\mathbb{G}(\mathbb{A}_{K})$ (cf. [5, Prop. 1.4.iv)]).
@@ 2.1 Number fields
In the remainder of this section, we will assume that $K$ is totally real. This means that $K$ has precisely $r=d_{K}$ distinct embeddings $\sigma_{1},\ldots,\sigma_{r}$ into ${\bf C}$ with $\sigma_{i}(K)\subseteq{\bf R}$ for all $i$. If $\mathcal{O}_{K}\subseteq K$ is the ring of algebraic integers of $K$, the multiplicative group $\mathcal{O}_{K}^{\times}$ of units is a finitely generated abelian group and can therefore be decomposed as $\mathcal{O}_{K}^{\times}\simeq\mu(K)\times U_{K}$, where $\mu(K)$ denotes the torsion subgroup, consisting of units of finite order, and $U_{K}$ denotes the free part. Then the statement of Dirichlet’s unit theorem [20, Thm. I.7.3] is that $U_{K}$ has rank $r-1=d_{K}-1$. Let us suppose that $K\neq{\bf Q}$ so that this rank is non-zero. Then, any set of $r-1$ units $\mathopen{}\mathclose{{}\left\{\varepsilon_{1},\ldots,\varepsilon_{r-1}}\right\}\subseteq\mathcal{O}_{K}^{\times}$ that generate the free group $U_{K}$ is called a fundamental system of units or a collection of fundamental units. For example, if $K$ is a quadratic number field, then a fundamental unit is given by
@@
Although the regulator of $K$ does not appear in relation to Prasad’s volume formula, it is closely related to an invariant of $K$ that appears in the crucial estimate (3.1) of the index $\mathopen{}\mathclose{{}\left[\Gamma:\Lambda}\right]$ of the normalizer $\Lambda$ of the lattice $\Gamma$ (see Section 2.3). This invariant is the class number $h_{K}$ of $K$ (see [20, §I.6]). The relation between $R_{K}$ and $h_{K}$ is given by the following version of the Brauer–Siegel formula, which appears at various places in the literature concerning the covolumes of arithmetic groups (e.g. [5, eq. (6.1)]).
@@
If we let $\textbf{H}=\mathopen{}\mathclose{{}\left\{\textbf{x}\in{\bf R}^{d_{K}}:N(\textbf{x})\geqslant 1}\right\}$ and $\lambda=\sqrt{D_{K}}\cdot\mathop{\mathrm{res}}_{s=1}\zeta_{K}(s)$, we obtain from [27, Lemma 1] that, for any $s\in{\bf C}$,
@@
By the class number formula [20, Corollary 5.11], we have $\lambda=2^{{d_{K}}-1}R_{K}h_{K}$. The claim now follows from this by substituting $s=t+1$ in (2.6), where $t>0$ is arbitrary. ∎
In order to handle the gamma function appearing in the statement of Proposition 2.1, we record the following ”pointwise” version of Stirling’s formula due to Robbins [25]. Thus, if $n\geqslant 1$ is arbitrary, one has the estimates
@@
where $S$ runs over all finite subsets of places containing all the infinite places. We recall that the restricted direct product defining $\mathbb{A}_{K}$ is characterized by the fact that, as far as the coordinates belonging to the finite places are concerned, at most finitely many coordinates lie outside of the ring $\mathcal{O}_{\upsilon}:=\mathcal{O}_{K_{\upsilon}}$ of integers in $K_{\upsilon}$. In particular, the number field $K$ embeds diagonally into $\mathbb{A}_{K}$. We refer to [22, Chapter 1] for more information.
@@ 2.2 Algebraic groups
We will now recall some standard definitions and results from the theory of linear algebraic groups (cf. also [22, Chapters 2 and 3]). Throughout this section, $K$ will denote a number field, and $\mathbb{G}$ will denote a linear algebraic group.
@@
The group $\mathrm{Sp}_{2n}({\bf R})$ has no outer forms. Indeed, from [22, Sect. 2.1.13] it follows in particular that the number of isomorphism classes of outer forms of a group is equal to the number of graph automorphisms of its Dynkin diagram. Since the diagram $C_{n}$ has no non-trivial automorphisms, our claim follows.
On the other hand, the (simply connected) inner forms of $\mathrm{Sp}_{2n}({\bf R})$ can be described in the following straightforward manner: By [22, Prop. 2.19] every simply connected inner form of a group of type $C_{n}$ is of the form $\mathrm{SU}_{m}(D,f)$ where $D$ is a central division algebra of index $2n/m\in{\bf Z}$ over $K$, equipped with an involutive antiautomorphism $\tau:D\rightarrow D$, and $f$ is a nondegenerate sesquilinear form, which is either Hermitian or skew-Hermitian (depending on the type of $\tau$ — see [22, Sect. 2]).
A subgroup $B\subseteq\mathbb{G}$ is called a Borel subgroup if it is both connected and solvable and not properly contained in any other connected, solvable subgroup of $\mathbb{G}$. $\mathbb{G}$ itself is called quasisplit over $K$ if it contains a Borel subgroup which is defined over $K$. An important fact is that over any field $K$, an algebraic group has a unique quasisplit inner form (cf. [6, Prop. 7.2.12]).
@@
where $K_{\upsilon}$ is the completion of $K$ at the infinite place $\upsilon$. For more details, we refer to [22, Sect. 2.1.2].
@@
Furthermore, as a consequence of the fact that $K$ embeds diagonally into $\mathbb{A}_{K}$, the same is true for the group $\mathbb{G}(K)$ of $K$-rational points, which in fact embeds as a discrete subgroup of $\mathbb{G}(\mathbb{A}_{K})$. We refer to [16, Sect. I.(0.33)] for more details.
@@
First of all, we see from [19, §18.5] that at least $K\subseteq{\bf R}$. Next, [19, Corollary 5.5.16] shows that $\mathbb{G}({\bf R})$ is simple, and that the diagonal embedding $\Delta(\mathbb{G}(\mathcal{O}_{K}))$ in $\mathrm{Res}_{K/{\bf Q}}(\mathbb{G})({\bf R})$ is therefore an irreducible lattice due to [19, Prop. 5.5.8 and Remark 5.5.9].
@@ 2.3.1 Parahoric subgroups
Just as one has the notions of Borel and parabolic subgroups of a reductive algebraic group, one has the notions of Iwahori and parahoric subgroups of an algebraic group over non-archimedean local fields. As a matter of fact, the following picture provides a useful intuition: if $K_{\upsilon}$ is the completion of a number field $K$ at a finite place $\upsilon$, and $\mathcal{O}_{\upsilon}\subseteq K_{\upsilon}$ is the valuation ring of $K_{\upsilon}$ containing the unique maximal ideal $\mathfrak{m}_{\upsilon}=\mathopen{}\mathclose{{}\left\{x\in\mathcal{O}_{\upsilon}:\upsilon(x)>0}\right\}$, one has the usual projection map $\mathcal{O}_{\upsilon}/\mathfrak{m}_{\upsilon}\twoheadrightarrow\mathbb{F}_{q_{\upsilon}}$ to the finite field with $q_{\upsilon}$ elements. An Iwahori subgroup $I_{\upsilon}\subseteq\mathbb{G}(K_{\upsilon})$ is then essentially the preimage $\pi^{-1}(B)$ of a Borel subgroup $B$ of $\overline{\mathbb{G}}$, the group $\mathbb{G}$ ”considered” as a group over the finite field $\mathbb{F}_{q_{\upsilon}}$. (This will be made precise below.) Furthermore, a parabolic subgroup of $\overline{\mathbb{G}}(\mathbb{F}_{q_{\upsilon}})$ is, by definition, any group containing (a conjugate of) $B$, and a parahoric subgroup $P_{\upsilon}\subseteq\mathbb{G}(K_{\upsilon})$ is then essentially the preimage $\pi^{-1}(P)$ of a parabolic subgroup.
@@
We recall the following facts from [22, Sect. 3.4]: There is an Iwahori subgroup $B_{\upsilon}$ in $\mathbb{G}(K_{\upsilon})$ and a maximal $K_{\upsilon}$-split torus $S_{\upsilon}\subseteq\mathbb{G}(K_{\upsilon})$ (say, of dimension $\dim S_{\upsilon}=\ell$) such that with $N_{\upsilon}=(N_{\mathbb{G}}(S_{\upsilon}))(K_{\upsilon})$ equal to the normalizer of $S_{\upsilon}$, the pair $(B_{\upsilon},N_{\upsilon})$ is a so-called $BN$-pair for the group of $K_{\upsilon}$-rational points $\mathbb{G}(K_{\upsilon})$. This means in particular that $H_{\upsilon}:=B_{\upsilon}\cap N_{\upsilon}$ is a normal subgroup in $N_{\upsilon}$, and that there is a size $\ell+1$ generating subset $\Delta_{\upsilon}=\mathopen{}\mathclose{{}\left\{r_{0},r_{1},\ldots,r_{\ell}}\right\}\subseteq W_{\upsilon}:=N_{\upsilon}/H_{\upsilon}$ of the Weyl group $W_{\upsilon}$, corresponding to the vertices in the local Dynkin diagram of $\mathbb{G}(K_{\upsilon})$, with every element of $\Delta_{\upsilon}$ having order $2$. It turns out that if $B_{\upsilon}\subseteq P_{\upsilon}$ for some subgroup $P_{\upsilon}\subseteq\mathbb{G}(K_{\upsilon})$, then there is a subset $\Theta_{\upsilon}\subseteq\Delta_{\upsilon}$ of generators, and a resulting subgroup $W_{\Theta_{\upsilon}}\subseteq W_{\Delta_{\upsilon}}=W_{\upsilon}$ generated by the elements of $\Theta_{\upsilon}$, such that $P_{\upsilon}=B_{\upsilon}W_{\Theta_{\upsilon}}B_{\upsilon}$. The subset $\Theta_{\upsilon}\subseteq\Delta_{\upsilon}$ is then called the type of $P_{\upsilon}$.
@@
We recall from [5, Sect. 3.1] that a parahoric subgroup $P_{\upsilon}$, which has maximal volume among all parahoric subgroups, is necessarily special; and from [5, Sect. 3.2] that a hyperspecial parahoric subgroup necessarily has maximal volume.
Since we want $\Gamma\subseteq\mathrm{Sp}_{2n}({\bf R})$ to be of minimal covolume, we are in fact free to assume that, at each finite place $\upsilon$, the parahoric subgroup $P_{\upsilon}\subseteq\mathbb{G}(K_{\upsilon})$ has maximal volume. Indeed, this follows immediately from [4, Sect. 4.3] (see also [5, Sect. 3.8]) which is valid for algebraic groups of all types. In view of the discussion above, the assumption that $\Gamma$ has minimal covolume then implies that each parahoric $P_{\upsilon}$ is special. In turn, this fact then implies that the type $\Theta_{\upsilon}$ of $P_{\upsilon}$ has no symmetries, as is shown in [4, Sect. 3.1].
Another convenient consequence of assuming that each parahoric is special is the following simple characterization of hyperspecial parahoric subgroups: for a finite place $\upsilon$,
@@
(Indeed, it follows immediately from the definition that $P_{\upsilon}$ is hyperspecial if $\mathbb{G}$ splits over $K_{\upsilon}$. Conversely, if $P_{\upsilon}\subseteq\mathbb{G}(K_{\upsilon})$ is hyperspecial, then $\mathbb{G}$ must be quasisplit over $K_{\upsilon}$ by [11, Prop. 10.2.1]. Accordingly, $\mathbb{G}$ splits over $K_{\upsilon}$ by uniqueness of the quasisplit inner form over any field.)
We now proceed to the discussion of parahoric subgroups in the context of Prasad’s volume formula. Additional details are given in [23, Sect. 2.2] and [11, Sect. 4.1].
@@
$e^{\prime}(P_{\upsilon})=q_{\upsilon}^{(\dim\textbf{M}_{\upsilon}-\dim\mathcal{M}_{\upsilon})/2}\frac{\#\mathcal{M}_{\upsilon}(\mathbb{F}_{q_{\upsilon}})}{\#\textbf{M}_{\upsilon}(\mathbb{F}_{q_{\upsilon}})}=e(P_{\upsilon})\frac{\#\mathcal{M}_{\upsilon}(\mathbb{F}_{q_{\upsilon}})}{q_{\upsilon}^{\dim\mathcal{M}_{\upsilon}}}.$ (2.10)
We note that $e^{\prime}(P_{\upsilon})$ is always a non-negative integer (cf. [24, Sect. 2.5]).
Of course, it is possible to simplify both of these expressions to some extent by using the explicit description of $\mathcal{M}_{\upsilon}$ as the standard split form of the symplectic group over the residue field $\mathbb{F}_{q_{\upsilon}}$. We will do this in the next section. For now, we note (cf. [24, Sect. 2.5]) that one always has the inequality
@@
For later use, we also note the following equivalences (cf. [23, Sect. 2.2]):

### v3

@@ 1 Introduction
By the celebrated arithmeticity theorem of Margulis (see e.g. [21, Thm. 16.3.1]), any irreducible lattice $\Lambda$ in a semisimple Lie group $G$ of real rank greater than $1$ is also an arithmetic subgroup of $G$. In particular, if such a Lie group $G$ is a simple, any lattice is arithmetic. That is, there exists a ${\bf Q}$-embedding $\iota:G\rightarrow\mathrm{GL}_{n}({\bf R})$ such that $\Lambda$ is commensurable to the group $G({\bf Z})=\iota^{-1}(\iota(G)\cap\mathrm{GL}_{n}({\bf Z}))$ of integer points of $G$. For our purposes, the utility of this identification of lattices with arithmetic subgroups is that arithmetic subgroups are, in some sense, easier to describe. However, as any individual arithmetic subgroup of $G$ is associated to a corresponding embedding of $G$ into $\mathrm{GL}_{n}({\bf R})$, investigating all arithmetic subgroups involves the abstraction of treating all ${\bf Q}$-isomorphic copies of $G$ in $\mathrm{GL}_{n}({\bf R})$ equally.
We therefore let $\mathbb{G}$ be an algebraic group over a number field $K$, for which there exists an embedding $\upsilon_{0}:K\rightarrow{\bf R}$ which induces an ${\bf R}$-isomorphism from $\mathrm{Sp}_{2n}$ (the standard split form of the symplectic group) to $\mathbb{G}$. This is enough to capture any candidate for $\Gamma$ that we will want to consider. Moreover, it leads to a useful alternative characterization of $\Gamma$: Rohlfs gave a cohomological criterion [28, Satz 3.5] (see also [6, Prop. 1.4]) for maximality of arithmetic subgroups, according to which the lattice $\Gamma\subseteq\mathrm{Sp}_{2n}({\bf R})$ must be conjugate to the normalizer of a so-called principal arithmetic subgroup $\Lambda$ of $\mathbb{G}(K)$. Such subgroups are distinguished by being topologically well-behaved in relation to the group $\mathbb{G}(\mathbb{A}_{K})$ of adelic points of $\mathbb{G}$. In particular, $\Lambda$ is given by the intersection $\mathbb{G}(K)\cap\Pi_{\upsilon<\infty}P_{\upsilon}$ for a certain (topologically coherent) family of compact-open (more precisely, parahoric) subgroups $P_{\upsilon}\subseteq\mathbb{G}(K_{\upsilon})$, where $\upsilon$ ranges over all finite places of $K$. These facts make the problem of computing the covolume of $\Lambda$, and thereby identifying a lattice $\Gamma$ of minimal covolume, amenable to a wide range of analytic and algebraic tools.
@@ 1.2 Notation and conventions
  • $\Lambda$ will denote the principal arithmetic subgroup of $\mathbb{G}_{\infty}$ defined by the collection of parahorics $\mathopen{}\mathclose{{}\left\{P_{\upsilon}:\upsilon<\infty}\right\}$, for which $\Gamma$ is its normalizer in the group $\mathbb{G}({\bf R})$ (cf. [6, Prop. 1.4.iv)]).
@@ 2.1 Number fields
In the remainder of this section, we will assume that $K$ is totally real. This means that $K$ has precisely $r=d_{K}$ distinct embeddings $\sigma_{1},\ldots,\sigma_{r}$ into ${\bf C}$ with $\sigma_{i}(K)\subseteq{\bf R}$ for all $i$. If $\mathcal{O}_{K}\subseteq K$ is the ring of algebraic integers of $K$, the multiplicative group $\mathcal{O}_{K}^{\times}$ of units is a finitely generated abelian group and can therefore be decomposed as $\mathcal{O}_{K}^{\times}\simeq\mu(K)\times U_{K}$, where $\mu(K)$ denotes the torsion subgroup, consisting of units of finite order, and $U_{K}$ denotes the free part. Then the statement of Dirichlet’s unit theorem [22, Thm. I.7.3] is that $U_{K}$ has rank $r-1=d_{K}-1$. Let us suppose that $K\neq{\bf Q}$ so that this rank is non-zero. Then, any set of $r-1$ units $\mathopen{}\mathclose{{}\left\{\varepsilon_{1},\ldots,\varepsilon_{r-1}}\right\}\subseteq\mathcal{O}_{K}^{\times}$ that generate the free group $U_{K}$ is called a fundamental system of units or a collection of fundamental units. For example, if $K$ is a quadratic number field, then a fundamental unit is given by
@@
Although the regulator of $K$ does not appear in relation to Prasad’s volume formula, it is closely related to an invariant of $K$ that appears in the crucial estimate (3.1) of the index $\mathopen{}\mathclose{{}\left[\Gamma:\Lambda}\right]$ (see Section 2.3). This invariant is the class number $h_{K}$ of $K$ (see [22, §I.6]). The relation between $R_{K}$ and $h_{K}$ is given by the following version of the Brauer–Siegel formula, which appears at various places in the literature concerning the covolumes of arithmetic groups (e.g. [6, eq. (6.1)]).
@@
If we let $\textbf{H}=\mathopen{}\mathclose{{}\left\{\textbf{x}\in{\bf R}^{d_{K}}:N(\textbf{x})\geqslant 1}\right\}$ and $\lambda=\sqrt{D_{K}}\cdot\mathop{\mathrm{res}}_{s=1}\zeta_{K}(s)$, we obtain from [29, Lemma 1] that, for any $s\in{\bf C}$,
@@
By the class number formula [22, Corollary 5.11], we have $\lambda=2^{{d_{K}}-1}R_{K}h_{K}$. The claim now follows from this by substituting $s=t+1$ in (2.6), where $t>0$ is arbitrary. ∎
In order to handle the gamma function appearing in the statement of Proposition 2.1, we record the following “pointwise” version of Stirling’s formula due to Robbins [27]. Thus, if $n\geqslant 1$ is arbitrary, one has the estimates
@@
where $S$ runs over all finite subsets of places containing all the infinite places. We recall that the restricted direct product defining $\mathbb{A}_{K}$ is characterized by the fact that, as far as the coordinates belonging to the finite places are concerned, at most finitely many coordinates lie outside of the ring $\mathcal{O}_{\upsilon}:=\mathcal{O}_{K_{\upsilon}}$ of integers in $K_{\upsilon}$. In particular, the number field $K$ embeds diagonally into $\mathbb{A}_{K}$. We refer to [24, Chapter 1] for more information.
@@ 2.2 Algebraic groups
We will now recall some standard definitions and results from the theory of linear algebraic groups (cf. also [24, Chapters 2 and 3]). Throughout this section, $K$ will denote a number field, and $\mathbb{G}$ will denote a linear algebraic group.
@@
The group $\mathrm{Sp}_{2n}({\bf R})$ has no outer forms. Indeed, from [24, Sect. 2.1.13] it follows in particular that the number of isomorphism classes of outer forms of a group is equal to the number of graph automorphisms of its Dynkin diagram. Since the diagram $C_{n}$ has no non-trivial automorphisms, our claim follows.
On the other hand, the (simply connected) inner forms of $\mathrm{Sp}_{2n}({\bf R})$ can be described in the following straightforward manner: By [24, Prop. 2.19] every simply connected inner form of a group of type $C_{n}$ is of the form $\mathrm{SU}_{m}(D,f)$ where $D$ is a central division algebra of index $2n/m\in{\bf Z}$ over $K$, equipped with an involutive antiautomorphism $\tau:D\rightarrow D$, and $f$ is a nondegenerate sesquilinear form, which is either Hermitian or skew-Hermitian (depending on the type of $\tau$ — see [24, Sect. 2]).
A subgroup $B\subseteq\mathbb{G}$ is called a Borel subgroup if it is both connected and solvable and not properly contained in any other connected, solvable subgroup of $\mathbb{G}$. $\mathbb{G}$ itself is called quasisplit over $K$ if it contains a Borel subgroup which is defined over $K$. An important fact is that over any field $K$, an algebraic group has a unique quasisplit inner form (cf. [7, Prop. 7.2.12]).
@@
where $K_{\upsilon}$ is the completion of $K$ at the infinite place $\upsilon$. For more details, we refer to [24, Sect. 2.1.2].
@@
Furthermore, as a consequence of the fact that $K$ embeds diagonally into $\mathbb{A}_{K}$, the same is true for the group $\mathbb{G}(K)$ of $K$-rational points, which in fact embeds as a discrete subgroup of $\mathbb{G}(\mathbb{A}_{K})$. We refer to [18, Sect. I.(0.33)] for more details.
@@
First of all, we see from [21, §18.5] that at least $K\subseteq{\bf R}$. Next, [21, Corollary 5.5.16] shows that $\mathbb{G}({\bf R})$ is simple, and that the diagonal embedding $\Delta(\mathbb{G}(\mathcal{O}_{K}))$ in $\mathrm{Res}_{K/{\bf Q}}(\mathbb{G})({\bf R})$ is therefore an irreducible lattice due to [21, Prop. 5.5.8 and Remark 5.5.9].
@@ 2.3.1 Parahoric subgroups
Just as one has the notions of Borel and parabolic subgroups of a reductive algebraic group, one has the notions of Iwahori and parahoric subgroups of an algebraic group over non-archimedean local fields.
As a matter of fact, the following picture provides a useful intuition: if $K_{\upsilon}$ is the completion of a number field $K$ at a finite place $\upsilon$, and $\mathcal{O}_{\upsilon}\subseteq K_{\upsilon}$ is the valuation ring of $K_{\upsilon}$ containing the unique maximal ideal $\mathfrak{m}_{\upsilon}=\mathopen{}\mathclose{{}\left\{x\in\mathcal{O}_{\upsilon}:\upsilon(x)>0}\right\}$, one has the usual projection map $\mathcal{O}_{\upsilon}/\mathfrak{m}_{\upsilon}\twoheadrightarrow\mathbb{F}_{q_{\upsilon}}$ to the finite field with $q_{\upsilon}$ elements. An Iwahori subgroup $I_{\upsilon}\subseteq\mathbb{G}(K_{\upsilon})$ is then essentially the preimage $\pi^{-1}(B)$ of a Borel subgroup $B$ of $\overline{\mathbb{G}}$, the group $\mathbb{G}$ “considered” as a group over the finite field $\mathbb{F}_{q_{\upsilon}}$. (This will be made precise below.) Furthermore, a parabolic subgroup of $\overline{\mathbb{G}}(\mathbb{F}_{q_{\upsilon}})$ is, by definition, any group containing (a conjugate of) $B$, and a parahoric subgroup $P_{\upsilon}\subseteq\mathbb{G}(K_{\upsilon})$ is then essentially the preimage $\pi^{-1}(P)$ of a parabolic subgroup.
@@
We recall the following facts from [24, Sect. 3.4]: There is an Iwahori subgroup $B_{\upsilon}$ in $\mathbb{G}(K_{\upsilon})$ and a maximal $K_{\upsilon}$-split torus $S_{\upsilon}\subseteq\mathbb{G}(K_{\upsilon})$ (say, of dimension $\dim S_{\upsilon}=\ell$) such that with $N_{\upsilon}=(N_{\mathbb{G}}(S_{\upsilon}))(K_{\upsilon})$ equal to the normalizer of $S_{\upsilon}$, the pair $(B_{\upsilon},N_{\upsilon})$ is a so-called $BN$-pair for the group of $K_{\upsilon}$-rational points $\mathbb{G}(K_{\upsilon})$. This means in particular that $H_{\upsilon}:=B_{\upsilon}\cap N_{\upsilon}$ is a normal subgroup in $N_{\upsilon}$, and that there is a size $\ell+1$ generating subset $\Delta_{\upsilon}=\mathopen{}\mathclose{{}\left\{r_{0},r_{1},\ldots,r_{\ell}}\right\}\subseteq W_{\upsilon}:=N_{\upsilon}/H_{\upsilon}$ of the Weyl group $W_{\upsilon}$, corresponding to the vertices in the local Dynkin diagram of $\mathbb{G}(K_{\upsilon})$, with every element of $\Delta_{\upsilon}$ having order $2$. It turns out that if $B_{\upsilon}\subseteq P_{\upsilon}$ for some subgroup $P_{\upsilon}\subseteq\mathbb{G}(K_{\upsilon})$, then there is a subset $\Theta_{\upsilon}\subseteq\Delta_{\upsilon}$ of generators, and a resulting subgroup $W_{\Theta_{\upsilon}}\subseteq W_{\Delta_{\upsilon}}=W_{\upsilon}$ generated by the elements of $\Theta_{\upsilon}$, such that $P_{\upsilon}=B_{\upsilon}W_{\Theta_{\upsilon}}B_{\upsilon}$. The subset $\Theta_{\upsilon}\subseteq\Delta_{\upsilon}$ is then called the type of $P_{\upsilon}$.
@@
Remarks.
1) We recall from [6, Sect. 3.1] that a parahoric subgroup $P_{\upsilon}$, which has maximal volume among all parahoric subgroups, is necessarily special; and from [6, Sect. 3.2] that a hyperspecial parahoric subgroup necessarily has maximal volume. (See also [6, Prop. A.5].)
2) Since we want $\Gamma\subseteq\mathrm{Sp}_{2n}({\bf R})$ to be of minimal covolume, we claim that we are, in fact, free to assume that at each finite place $\upsilon$, the parahoric subgroup $P_{\upsilon}\subseteq\mathbb{G}(K_{\upsilon})$ has maximal volume. Namely, this follows from the following two points: First of all, the argument in [5, Sect. 4.3] (which builds on [6, Sect. 3.8]) shows that the covolume of $\Gamma=N_{\mathbb{G}({\bf R})}(\Lambda)$ can only decrease when a parahoric subgroup $P_{\upsilon}$ in the coherent sequence of $\Lambda$ is replaced with a parahoric of larger volume, in case $P_{\upsilon}$ is not already maximal. (Although the statements made in [5] pertain to groups of type $D_{n}$, the proof of the inequality after [5, eq. 4.(15)] extends almost immediately to groups of type $C_{n}$ as well, with only minor adjustments required.) Second, we will show that the covolume must, in fact, decrease in this situation. This argument will rely on the analysis that follows (where every $P_{\upsilon}$ is assumed to be maximal) and, in particular, on Theorem 2.2 below, which we have not yet discussed. For this reason, we postpone the justification of this claim to Section 4 where the result is stated in the form of Lemma 4.3.
Assume from now on that each $P_{\upsilon}$ is special. This fact has at least the following useful consequences. First of all, the type $\Theta_{\upsilon}$ of any parahoric $P_{\upsilon}$ in our sequence has no symmetries, as is shown in [8, Sect. 3.1]. Second, the following simple characterization of hyperspecial parahoric subgroups holds: for a finite place $\upsilon$,
@@
Indeed, it follows immediately from the definition that $P_{\upsilon}$ is hyperspecial if $\mathbb{G}$ splits over $K_{\upsilon}$. Conversely, if $P_{\upsilon}\subseteq\mathbb{G}(K_{\upsilon})$ is hyperspecial, then $\mathbb{G}$ must be quasisplit over $K_{\upsilon}$ by [13, Prop. 10.2.1]. Accordingly, $\mathbb{G}$ splits over $K_{\upsilon}$ by uniqueness of the quasisplit inner form over any field.
We now proceed to the discussion of parahoric subgroups in the context of Prasad’s volume formula. Additional details are given in [25, Sect. 2.2] and [13, Sect. 4.1].
@@
$e^{\prime}(P_{\upsilon}):=q_{\upsilon}^{(\dim\textbf{M}_{\upsilon}-\dim\mathcal{M}_{\upsilon})/2}\frac{\#\mathcal{M}_{\upsilon}(\mathbb{F}_{q_{\upsilon}})}{\#\textbf{M}_{\upsilon}(\mathbb{F}_{q_{\upsilon}})}=e(P_{\upsilon})\frac{\#\mathcal{M}_{\upsilon}(\mathbb{F}_{q_{\upsilon}})}{q_{\upsilon}^{\dim\mathcal{M}_{\upsilon}}}.$ (2.10)
We note that $e^{\prime}(P_{\upsilon})$ is always a non-negative integer (cf. [26, Sect. 2.5]).
Of course, it is possible to simplify both of these expressions to some extent by using the explicit description of $\mathcal{M}_{\upsilon}$ as the standard split form of the symplectic group over the residue field $\mathbb{F}_{q_{\upsilon}}$. We will do this in the next section. For now, we note (cf. [26, Sect. 2.5]) that one always has the inequality
@@
For later use, we also note the following equivalences (cf. [25, Sect. 2.2]):